Prove that


(a) (sin x + cos x) (tan x + cot x) = sec x + cosec x.


(b) sec x sin x / tan x + cot x = sin² x .

1
Expert's answer
2021-11-08T17:29:39-0500

Explanations & Calculations

"\qquad\qquad
\begin{aligned}
\small (\sin x + \cos x)(\tan x+\cot x)&=\small (\sin x+\cos x)\Big(\frac{\sin x}{\cos x}+\frac{\cos x}{\sin x}\Big)\\
&=\small (\sin x+\cos x)\Big(\frac{\sin^2 x+\cos^2 x}{\sin x \cos x}\Big)\\
&=\small \frac{1}{\cos x}+\frac{1}{\sin x}\\
&=\small \sec x+ \cosec x\\
\\
\\
\small \frac{\sec x \sin x}{\tan x+ \cot x}&=\small \frac{\tan x}{\tan x+\large\frac{1}{\tan x}}\\
&=\small \frac{\tan^2 x}{\tan^2 x+1}\\
&=\small \frac{\tan^2 x}{\sec^2 x}\\
&=\small \frac{\sin^2x}{\cos^2x\times\sec^2x}\\
&=\small \sin^2x
\end{aligned}"